jack is 23 years younger than his father.his father is x years old and the sum of their ages is 64.calculate jack's age
Answer by lwsshak3(11628)

jack is 23 years younger than his father.his father is x years old and the sum of their ages is 64.calculate jack's age
***
let x=father's age
x-23=jack's age
x+(x-23)=64
2x-23=64
2x=87
x=43.5
x-23=20.5
jack's age=20.5